The sum of the interior angles of a polygon is 540°. How many sides does this polygon have?
Answer by gwendolyn(128)

The sum of the interior angles of a polygon is equal to 2 less than the number of sides the polygon has multiplied by 180.
This can be expressed in an equation as:
a=180(s-2) where a equals the sum of interior angles and s equals the number of sides the polygon has.
If we plug 540 in as the 'a' value, we find that:
540=180(s-2)
We can then divide both sides of the equation by 180:
3=s-2
Finally, we can add 2 to both sides of the problem:
5=s
Therefore we find that the polygon has 5 sides.